OpenStudy (anonymous):

Two similar cylinders have heights 12 inches and 18 inches. What is the ratio of their surface areas?

OpenStudy (anonymous):

@Ganpat @SmoothMath @experimentX

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (ganpat):

Surface are for cylinder.. Surface Area = Areas of top and bottom +Area of the side Surface Area = 2(Area of top) + (perimeter of top)* height since are is not given... i will assume the radius will be the same.. So A1/ A2 = (perimeter of top)1 * height 1 / (perimeter of top)2 * height22 A1 / A2 = 2*pi*r* h1 / 2* pi*r* h2.. A1 / A2 = 12 / 18 = 2 /3...

OpenStudy (anonymous):

Ok. The ratio between two similar figures' sides and lengths is a : b. The ratio between areas is \(a^2 : b^2\).So if a : b is 12 : 18 which when simplifies is 2 : 3, what is \(a^2 : b^2\)?

OpenStudy (anonymous):

THANKS GUYS

OpenStudy (anonymous):

np :)